5.8 liters of oxygen, O2, to molecules of oxygen
1
Expert's answer
2020-05-15T09:36:57-0400

N(O2)=(V(O2)/VM)*NA=(5.8/22.4)*6.02*10^23=1.56 * 10^23 molecules of Oxygen